OpenStudy (anonymous):

How about this one. (6-sqrt(6))(5+sqrt(6)). The instructions say to multiply and simplify answer. Type and exact answer, using radicals as needed.

OpenStudy (anonymous):

This is what it looks like.\[(6-\sqrt{6})(5+\sqrt{6})\]

OpenStudy (anonymous):

use the foil method -

OpenStudy (anonymous):

foil....got it...let me give it a try.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got 26.449

OpenStudy (anonymous):

ok nope...trying that again.

OpenStudy (saifoo.khan):

Your answer's correct though. But in online test, it isnt going to work! D:

OpenStudy (anonymous):

ok so foil...first multiply 6*5=35, then multiply 6*sqrt(6), that is were i get lost. How do you multiply that?

OpenStudy (anonymous):

that is it 6 srt 6

OpenStudy (anonymous):